Show that 2°is equal to 1


1
Expert's answer
2021-02-12T06:13:01-0500

We all know the rule:


"\frac{a^x}{a^y}=a^{x-y}."

Consider this:


"\frac{2^x}{2^y}=2^{x-y}."

Now, consider this:


"\frac{2^x}{2^x}=2^{x-x}=2^0."


But what is "2^x\/2^x"?

According to the definition of power x:

"2^x=2·2·2·...·2·2\text{ (the number of 2 used is }\textit{x}\text{)}."


So, since we take the same number of 2s and divide it by the equal number of 2, we divide a value by itself, which gives 1. Thus, "2^0=1".